OpenStudy (anonymous):

Change this radical to an algebraic expression with fractional exponents. √y^2= ... Please help!

OpenStudy (aravindg):

what do you know about square root of a number ?

OpenStudy (anonymous):

Not very much. I just need help with this problem very badly.

OpenStudy (aravindg):

if \[\large \bf x^2 =k \] then \[\large \bf x=\pm \sqrt{k}\]

OpenStudy (aravindg):

\[\large \bf +\sqrt{k}\] is known as positive square root or simply square root of x

OpenStudy (aravindg):

thus we can see than 5 will be square root of 25, 4 will be square root of 16 etc

OpenStudy (aravindg):

now can you do your question ?

OpenStudy (anonymous):

Well the question wasn't about square roots. It was about radicals, so no, I still can't.

OpenStudy (aravindg):

of course it was related to square root ! the question asks for \[\large \bf \sqrt{y^2} \]

OpenStudy (aravindg):

so find square root of y^2